A gas is compressed from 120 kPa and 15oC to 800 kPa. Calculate the final temperature when

the process is

i. Isothermal (n=1)

ii. Polytropic (n=1.3)

iii Adiabatic (γ=1.4)

iv. Polytropic (n= 1.5)


1
Expert's answer
2021-09-04T18:21:12-0400

Solution.

i. "T_2=T_1=15^oC;"

ii. "T2 = T1 (p2 \/ p1)^{\n1-1\/n} = 288(800\/120)^{0.231} = 446 K= 173^oC;"

III."T_2 = T_1 (p_2 \/ p_1)^{\n1-1\/\u03b3}\n = 288(800\/120)^{0.286} = 495 K= 222^oC;"

iv."T_2 = T_1 (p_2 \/ p_1)^{\n1-1\/n }= 288(800\/120)^{0.333} = 541 K =269^oC;"

Answer: "i. T_2=15^oC;"

"ii. T_2=173^oC;"

"iii. T_2=222^oC;"

"iv. T_2=269^oC;"